1. In a large population of high school students, 20% have experienced math anxiety. You take a random sample of 10 students from this population. The standard deviation of the number of students in the sample who have experienced math anxiety is:

a. 0.0160

b. 0.2070

c. 0.2530

d. 1.0000

e. 1.2650

2. In a certain large population, 40% of households have a total annual income of $70,000. A simple random sample of 4 of these households is randomly selected. What is the probability that 2 or more of the households in the survey have an annual income of over $70,000?

a. 0.5248

b. 0.5000

c. 0.4000

d. 0.3456

e. 0.1600

3. A factory makes processors for use in computers. It is known that about 90% of the processors meet specifications. Every week a sample of 18 processors is selected at random for testing. Assume a binomial distribution is valid. Suppose we collect a large number of these samples of 18 processors and determine the number meeting specifications in each sample. What is the mean number of processors meeting specifications?

a. 1.62

b. 16.2

c. 4.02

d. 16

e. 22.3
Answer #1

Question 1

Mean = n * P = ( 10 * 0.2 ) = 2
Varaince = n * P * Q = ( 10 * 0.2 * 0.8 ) = 1.6
Standard deviation = = 1.2650

Option e is correct.

Question 2






P ( X >= 2 ) = 0.5248

Option A is correct.

Question 3

Option B is correct.